OpenStudy (anonymous):

find the derivitive of the lim as x approaches pi/2 for sin(x)-1/sqrt(sinx) -1

OpenStudy (anonymous):

L'Hospital rule

OpenStudy (jmark):

\[Given \lim x->\pi/2 for \sin(x)-1/\sqrt(sinx) -1\] \[\lim x->\pi/2 for [(√sinx -1)(√sinx+1)]/(√(sinx) -1)\] \[\lim x->\pi/2 for √sinx +1\] apply limit \[ √\sin(\pi/2)+1\] 1+1 2 ICSE Question Papers Class 9 - http://education-help-online.blogspot.in/2014/09/icse-question-papers-class-9.html